# Bayesian Group Regularization in Generalized Linear Models with a Continuous Spike-and-Slab Prior thanks: Keywords and phrases: generalized linear model, group sparsity, spike-and-slab, spike-and-slab group lasso, posterior contraction, rate of convergence, variable selection

### v7

@@ C.2 M-step for updating the regression coefficients
Recall that $\bm{\beta}=(\bm{\beta}_{1}^{\top},\ldots,\bm{\beta}_{G}^{\top})^{\top}$, where $\bm{\beta}_{g}\in\mathbb{R}^{m_{g}},g=1,\ldots,G$, and $\bm{\tau}=(\tau_{1},\ldots,\tau_{G})^{\top}$ is defined in (5.14). Let $\mathcal{PG}(a,b)$ denote the Pólya-gamma distribution [31], and let $\mathcal{GIG}(a,b,c)$ denote the generalized inverse Gaussian distribution with density function $f(x;a,b,c)\propto x^{c-1}e^{-(a/x+bx)/2}$. We define $\bm{\kappa}=(\kappa_{1},\ldots,\kappa_{n})^{\top}$, where for $i=1,\ldots,n$,
@@
The Gibbs sampler for grouped logistic or Poisson regression with the SSGL prior (5.14) is given in Algorithm 1.
@@
  •       i.   Sample $\gamma_{g}^{(t)}\sim\text{Bernoulli}\left(\frac{\pi_{1}}{\pi_{1}+\pi_{0}}\right)$, where $\pi_{1}=\theta^{(t-1)}\lambda_{1}^{m_{g}+1}\exp\left(-\lambda_{1}^{2}\tau_{g}^{(t-1)}/2\right)$          and $\pi_{0}=\left(1-\theta^{(t-1)}\right)\lambda_{0}^{m_{g}+1}\exp\left(-\lambda_{0}^{2}\tau_{g}^{(t-1)}/2\right)$       ii. Sample $\tau_{g}^{(t)}\sim\mathcal{GIG}\left(\frac{1}{2},\lVert\bm{\beta}_{g}^{(t-1)}\rVert_{2}^{2},(\lambda_{g}^{\star})^{2}\right)$, where $\lambda_{g}^{\star}=\tau_{g}^{(t)}\lambda_{1}+(1-\tau_{g}^{(t)})\lambda_{0}$
@@
    When $p=\sum_{g=1}^{G}m_{g}$ is larger than $n$, we can use the fast sampling algorithm of Bhattacharya et al. [5] to sample $\bm{\beta}^{(t)}$ in Step 4 of Algorithm 1 in $\mathcal{O}(n^{2}p)$ time instead of $\mathcal{O}(p^{3})$ time. This sampling algorithm is given in Algorithm 2.
@@ Appendix E Simulation studies for negative binomial regression with a log link
    We now present simulation results in grouped negative binomial regression for the SSGL model (2.2). We compared our results to group lasso (GL), adaptive group lasso (AdGL), group SCAD (GSCAD), and group MCP (gMCP). We used the same methods for choosing the hyperparameters, regularization parameters, and weights as those described in Section 6.1. For our simulation study, we fixed $\alpha=1$ and generated the responses independently as $y_{i}\mid\mathbf{x}_{i}\sim NB(\alpha,\exp(\theta_{i})),i=1,\ldots,n$. We conducted the following simulation studies:
@@
    We repeated each experiment 200 times. Table E1 reports our results averaged across the 200 replications. In Experiment 1, SSGL had the lowest MSE and the second lowest MSPE. Meanwhile, all five methods had a TPR of 1, indicating that all methods had equally high power to detect the significant groups in this particular setting. However, SSGL had the highest average TNR and precision, indicating superior overall group selection performance.

### v8

@@ B.2 Proof of Theorem 3
B.3 Proof of Proposition 1
Throughout, we use $P_{0}$ and $\mathbb{E}_{0}$ to respectively denote the probability and expectation operators under the law induced by the model $\mathbf{Y}\sim\mathcal{N}_{n}(\bm{\beta}_{0},\mathbf{I}_{n})$.
The posterior contraction proof in Proposition 1 follows a similar strategy as the proof of Theorem 7 in Castillo et al. [11]. However, it should be noted that Castillo et al. [11] deal exclusively with the case where $\bm{\beta}_{0}=(\beta_{01},\ldots,\beta_{0n})^{\top}$ is the zero vector, i.e. $\bm{\beta}_{0}=\bm{0}_{n}$, and the individual regression coefficients $\beta_{0i},i=1,\ldots,n$, are endowed with univariate $\text{Laplace}(0,\lambda^{-1})$ priors. In contrast, we consider the grouped scenario where $\bm{\beta}_{0}=(\bm{\beta}_{01}^{\top},\ldots,\bm{\beta}_{0G}^{\top})^{\top}$ with $\bm{\beta}_{01}\neq\bm{0}_{m_{1}}$, and the groups $\bm{\beta}_{0g},g=1,\ldots,G$, are endowed with independent multivariate Laplace priors (2.1).
Proof of Proposition 1
We first prove (4.6). Let $\mathbf{Y}=(\mathbf{Y}_{1}^{\top},\ldots,\mathbf{Y}_{G}^{\top})^{\top}$, where the indices of $\mathbf{Y}_{g}$ correspond to those of $\bm{\beta}_{0g}$ in $\bm{\beta}_{0}=(\bm{\beta}_{01}^{\top},\ldots,\bm{\beta}_{0G}^{\top})^{\top}$ for all $g=1,\ldots,G$. When $\mathbf{Y}\sim\mathcal{N}_{n}(\bm{\beta}_{0},\mathbf{I}_{n})$, the group lasso estimator has a closed form,
$\widehat{\bm{\beta}}_{g}=\left(1-\frac{\lambda}{\lVert\mathbf{Y}_{g}\rVert_{2}}\right)_{+}\mathbf{Y}_{g},~{}\text{ for all }~{}g=1,\ldots,G,$
where $x_{+}=\min\{x,0\}$. Therefore, we must have $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}<\lVert\mathbf{Y}_{g}\rVert_{2}$ for all $g\in\{1,\ldots,G\}$. We can decompose the expected squared $\ell_{2}$ risk as
$\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}-\bm{\beta}_{0}\rVert_{2}^{2}=\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}_{1}-\bm{\beta}_{01}\rVert_{2}^{2}+\sum_{g=2}^{G}\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{2},$ (B.27)
We first upper-bound the first term in (B.27). Since $\mathbf{Y}_{1}\sim\mathcal{N}_{m}(\bm{\beta}_{01},\mathbf{I}_{m})$, $\lVert\widehat{\bm{\beta}}_{1}\rVert_{2}<\lVert\mathbf{Y}_{1}\rVert_{2}$, and $\widehat{\bm{\beta}}_{1}=\bm{0}_{m}$ when $\lVert\mathbf{Y}_{1}\rVert_{2}\leq\lambda$, we have
$\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}_{1}-\bm{\beta}_{01}\rVert_{2}^{2}$ $\leq 2\mathbb{E}_{0}\lVert\mathbf{Y}_{1}-\bm{\beta}_{01}\rVert_{2}^{2}+2\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}_{1}-\mathbf{Y}_{1}\rVert_{2}^{2}$
$\leq 2m+2\mathbb{E}_{0}\{\lVert\mathbf{Y}_{1}\rVert_{2}^{2}\mathbb{I}(\lVert\mathbf{Y}_{1}\rVert_{2}\leq\lambda)\}+2\mathbb{E}_{0}\{\lambda^{2}\mathbb{I}(\lVert\mathbf{Y}_{1}\rVert_{2}>\lambda)\}$
$\leq 2m+2\lambda^{2}+2\lambda^{2}P_{0}(\lVert\mathbf{Y}_{1}\rVert_{2}>\lambda)$
$\leq 2m+4\lambda^{2}\asymp 2\log n,$ (B.28)
since $m$ is constant and $\lambda\asymp\sqrt{2\log n}$. Next, we upper-bound the second term in (B.27). Again, using the fact that $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}<\lVert\mathbf{Y}_{g}\rVert_{2}$ and $\widehat{\bm{\beta}}_{g}=\bm{0}_{m}$ when $\lVert\mathbf{Y}_{g}\rVert_{2}\leq\lambda$, each of the summands can be upper-bounded as
$\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{2}$ $\leq\mathbb{E}_{0}\{\lVert\mathbf{Y}_{g}\rVert_{2}^{2}\mathbb{I}(\lVert\mathbf{Y}_{g}\rVert_{2}^{2}>\lambda^{2})\}.$
[…]
$=2\left[\Gamma(\frac{m}{2})\right]^{-1}\Gamma\left(\frac{m}{2}+1,\lambda^{2}\right)$
$\leq 2\left[\Gamma\left(\frac{m}{2}\right)\right]^{-1}\left(\frac{m}{2}+1\right)e^{-\lambda^{2}}(\lambda^{2})^{m/2}~{}\text{ for large }n$
$\asymp n^{-2}(\log n)^{m/2},$ (B.29)
where we used the bound $\Gamma(a,x)\leq ae^{-x}x^{a-1}$ when $a\geq 1$ and $x>a$ [32] in the third line, and the fact that $m$ is fixed and $\lambda\asymp\sqrt{2\log n}$ in the last line. Combining (B.27)-(B.3) gives
$\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}-\bm{\beta}_{0}\rVert_{2}^{2}\lesssim 2\log n+(G-1)n^{-2}(\log n)^{m/2}=2\log n+o(1)~{}\text{ as }n\rightarrow\infty,$
where we used the fact that $G=n/m$ and $m$ is a constant. This completes the proof of (4.6).
Next, we prove (4.7). We follow the proof strategy of Theorem 7 in Castillo et al. [11]. According to Lemma 7.1 of [12], $\mathbb{E}_{0}\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<s_{n}\mid\mathbf{Y})\rightarrow 0$ for any $s_{n}$ for which there exists a sequence $r_{n}$ such that
$\frac{\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<s_{n})}{\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<r_{n})}e^{2r_{n}^{2}}=o(1).$ (B.30)
Since the prior is $\bm{\beta}_{g}\overset{iid}{\sim}\Psi(\bm{\beta}_{g}\mid\lambda)$ for all $g=1,\ldots,G$, $\lVert\bm{\beta}_{1}\rVert_{2}\sim\text{Gamma}(m,\lambda)$ and $\sum_{g=1}^{G}\lVert\bm{\beta}_{g}\rVert_{2}\sim\text{Gamma}(n,\lambda)$ where $n=mG$.
We first upper-bound the numerator in (B.30). Since $\lVert\bm{\beta}_{0}\rVert_{2}=\lVert\bm{\beta}_{01}\rVert_{2}=L$ and $\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\geq\lVert\bm{\beta}_{1}-\bm{\beta}_{01}\rVert_{2}$, we have that for sufficiently large $n$ such that $s_{n}>L$,
$\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<s_{n})$ $\leq\Pi(\bm{\beta}_{1}:\lVert\bm{\beta}_{1}-\bm{\beta}_{0}\rVert_{2}<s_{n})$
$\leq\Pi\left(\bm{\beta}_{1}:\big{|}\lVert\bm{\beta}_{1}\rVert_{2}-\lVert\bm{\beta}_{01}\rVert_{2}\big{|}<s_{n}\right)$
$=\Pi(\bm{\beta}_{1}:\lVert\bm{\beta}_{1}\rVert_{2}<s_{n}+L)$
[…]
$\leq\frac{[\lambda(s_{n}+L)]^{m}}{\Gamma(m+1)}.$ (B.31)
Next, we lower-bound the denominator in (B.30). For sufficiently large $n$ so that $r_{n}>L$, we have
$\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<r_{n})$ $\geq\Pi(\bm{\beta}:\lVert\bm{\beta}\rVert_{2}+\lVert\bm{\beta}_{0}\rVert_{2}<r_{n})$
$=\Pi(\bm{\beta}:\lVert\bm{\beta}\rVert_{2}<r_{n}-L)$
$\geq\Pi\left(\bm{\beta}:\sum_{g=1}^{G}\lVert\bm{\beta}_{g}\rVert_{2}<r_{n}-L\right)$
[…]
$\geq\frac{\lambda^{n}}{n\Gamma(n)}e^{-\lambda(r_{n}-L)}(r_{n}-L)^{n}.$ (B.32)
Combining (B.30)-(B.3) gives
$\frac{\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<s_{n})}{\Pi(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<r_{n})}$ $\leq n\lambda^{m-n}\frac{\Gamma(n)}{\Gamma(m+1)}\frac{(s_{n}+L)^{m}}{(r_{n}-L)^{n}}e^{\lambda(r_{n}-L)}$
$\lesssim e^{\lambda\sqrt{n}r_{n}}\left(\frac{s_{n}}{r_{n}}\right)^{n}$
$=\exp\left[\lambda\sqrt{n}r_{n}-n\log\left(\frac{r_{n}}{s_{n}}\right)\right].$ (B.33)
Setting $r_{n}=\sqrt{n}(\lambda^{-1}\wedge 1)$ and $s_{n}=\delta r_{n}$ for $\delta$ small enough in (B.3) satisfies (B.30), and thus, by Lemma 7.1 of Castillo and van der Vaart [12],
$\mathbb{E}_{0}\Pi\left(\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}<\delta\sqrt{n}\left(\frac{1}{\lambda}\wedge 1\right)~{}\bigg{|}~{}\mathbf{Y}\right)\rightarrow 0~{}\textrm{ as }~{}n\rightarrow\infty.$
By setting $\lambda\asymp\sqrt{2\log n}$ in the above display, we obtain the posterior contraction rate of $\sqrt{n/\log n}$ in (4.7). $\square$
@@ C.2 M-step for updating the regression coefficients
The SSGL MAP algorithm is on average slower than the numerical algorithms used for group lasso, group SCAD, or group MCP. This is because each iteration of the EM algorithm requires numerically solve the optimization problem (C.3) in the M-step. Therefore, if the EM algorithm takes $B$ total iterations to converge, we expect the average runtime of the SSGL method to be on the order of $B$ times slower than that for group lasso, group SCAD, and group MCP. However, we observed that our EM algorithm tended to converge within a small number of iterations (typically fewer than 10), at least in all of the simulation settings we considered. Thus, the computation time for SSGL for fixed $\lambda_{0}$ was still reasonably quick. For example, in Simulations 4 and 8 of Section 6 where $p=1600$, the SSGL runtime for any fixed $\lambda_{0}$ was under 30 seconds in all replications.
To further accelerate the computational efficiency for SSGL, we could employ alternative optimization strategies that obviate the need to introduce latent indicator variables. For example, we could potentially approximate the SSGL prior with a neuronized prior [39], which is a product of a Gaussian weight variable and a scale variable transformed from Gaussian via an activation function. This approximation would allow for a fast coordinate ascent algorithm which avoids the need to marginalize out latent indicator variables [39]. We leave the investigation of more computationally efficient algorithms for SSGL for future work.
@@
Recall that $\bm{\beta}=(\bm{\beta}_{1}^{\top},\ldots,\bm{\beta}_{G}^{\top})^{\top}$, where $\bm{\beta}_{g}\in\mathbb{R}^{m_{g}},g=1,\ldots,G$, and $\bm{\tau}=(\tau_{1},\ldots,\tau_{G})^{\top}$ is defined in (5.14). Let $\mathcal{PG}(a,b)$ denote the Pólya-gamma distribution [33], and let $\mathcal{GIG}(a,b,c)$ denote the generalized inverse Gaussian distribution with density function $f(x;a,b,c)\propto{\color[rgb]{0,0,0}x^{a-1}e^{-(b/x+cx)/2}}$. We define $\bm{\kappa}=(\kappa_{1},\ldots,\kappa_{n})^{\top}$, where for $i=1,\ldots,n$,
@@
The Gibbs sampler for grouped logistic or Poisson regression with the SSGL prior (5.4) is given in Algorithm 1. When $p=\sum_{g=1}^{G}m_{g}$ is larger than $n$, we can use the fast sampling algorithm of Bhattacharya et al. [5] to sample $\bm{\beta}^{(t)}$ in Step 4 of Algorithm 1 in $\mathcal{O}(n^{2}p)$ time instead of $\mathcal{O}(p^{3})$ time. This sampling algorithm is given in Algorithm 2.
@@
  •       i.   Sample $\gamma_{g}^{(t)}\sim\text{Bernoulli}\left(\frac{\pi_{1}}{\pi_{1}+\pi_{0}}\right)$, where $\pi_{1}=\theta^{(t-1)}\lambda_{1}^{m_{g}+1}\exp\left(-\lambda_{1}^{2}\tau_{g}^{(t-1)}/2\right)$          and $\pi_{0}=\left(1-\theta^{(t-1)}\right)\lambda_{0}^{m_{g}+1}\exp\left(-\lambda_{0}^{2}\tau_{g}^{(t-1)}/2\right)$       ii. Sample $\tau_{g}^{(t)}\sim\mathcal{GIG}\left(\frac{1}{2},\lVert\bm{\beta}_{g}^{(t-1)}\rVert_{2}^{2},(\lambda_{g}^{\star})^{2}\right)$, where $\lambda_{g}^{\star}=\gamma_{g}^{(t)}\lambda_{1}+(1-\gamma_{g}^{(t)})\lambda_{0}$
@@ Appendix E Simulation studies for negative binomial regression with a log link
    We now present simulation results in grouped negative binomial regression for the SSGL model (2.2). We compared our results to group lasso (GL), adaptive group lasso (AdGL), group SCAD (GSCAD), and group MCP (gMCP). We used the same methods for choosing the hyperparameters, regularization parameters, and weights as those described in Section 6.1. We fixed $\alpha=1$ and generated the responses independently as $y_{i}\mid\mathbf{x}_{i}\sim NB(\alpha,\exp(\theta_{i})),i=1,\ldots,n$. We conducted the following simulation studies:
@@
    We repeated each experiment 200 times. Table 3 reports our results averaged across the 200 replications. In Experiment 1, SSGL had the lowest MSE and the second lowest MSPE. Meanwhile, all five methods had a TPR of 1, indicating that all methods had equally high power to detect the significant groups in this particular setting. However, SSGL had the highest average TNR and precision, indicating superior overall group selection performance.